OpenStudy (anonymous):

Find the value of Y for a given value of X, if Y varies directly with X. If Y =2.52 when X = 8.4, what is Y when X= 2.7? A. 9 B. -9 C. 0.81 D. -0.81

OpenStudy (anonymous):

Is it a proportion?

OpenStudy (cj49):

y=kx 2.52=k8.4 k=2.52/8.4 k=0.3 y=03x2.7 y=0.81

OpenStudy (anonymous):

Thank you so much! :) at was super easy for you huh? -_- I feel dorky...thanks dude.

OpenStudy (cj49):

lol...np...y varies directly with x means y=kx ( k is the constant)